One year​ ago, your company purchased a machine used in manufacturing for $115,000. You have learned that a new machine is available that offers many​ advantages; you can purchase it for $140,000 today. It will be depreciated on a​ straight-line basis over ten​ years, after which it has no salvage value. You expect that the new machine will contribute EBITDA​ (earnings before​ interest, taxes,​ depreciation, and​ amortization) of $60,000 per year for the next ten years. The current machine is expected to produce EBITDA of $22,000 per year. The current machine is being depreciated on a​ straight-line basis over a useful life of 11​ years, after which it will have no salvage​ value, so depreciation expense for the current machine is $10,455 per year. All other expenses of the two machines are identical. The market value today of the current machine is $50,000. Your​ company's tax rate is 45%​, and the opportunity cost of capital for this type of equipment is 10%. Is it profitable to replace the​ year-old machine?

The NPV of the replacement is ​$____. ​(Round to the nearest​ dollar.)
Answer #1

current book value of old machine = purchase price-1 year depreciation = 115000-10455=104545

incremental EBIDTA = EBIDTA new machine-EBIDTA old machine = 60000-22000=38000

Time line 0 1 2 3 4 5 6 7 8 9 10
Proceeds from sale of existing asset =selling price* ( 1 -tax rate) 27500
Tax shield on existing asset book value =Book value * tax rate 47045.25
Cost of new machine 140000
=Initial Investment outlay 214545.25
Incremental EBIDTA 38000 38000 38000 38000 38000 38000 38000 38000 38000 38000
-Depreciation Cost of equipment/no. of years 14000 14000 14000 14000 14000 14000 14000 14000 14000 14000
=Pretax cash flows 52000 52000 52000 52000 52000 52000 52000 52000 52000 52000
-taxes =(Pretax cash flows)*(1-tax) 28600 28600 28600 28600 28600 28600 28600 28600 28600 28600
+Depreciation -14000 -14000 -14000 -14000 -14000 -14000 -14000 -14000 -14000 -14000
=after tax operating cash flow 14600 14600 14600 14600 14600 14600 14600 14600 14600 14600
+Tax shield on salvage book value =Salvage value * tax rate 0
=Terminal year after tax cash flows 0
Total Cash flow for the period 214545.25 14600 14600 14600 14600 14600 14600 14600 14600 14600 14600
Discount factor= (1+discount rate)^corresponding period 1 1.1 1.21 1.331 1.4641 1.61051 1.771561 1.948717 2.14359 2.357948 2.593742
Discounted CF= Cashflow/discount factor 214545.25 13272.72727 12066.1157 10969.196 9971.9964 9065.4513 8241.3194 7492.109 6811.01 6191.825 5628.932
NPV= Sum of discounted CF= 304255.93

NPV is positive hence it is profitable